OpenStudy (anonymous):

Solve the following system. (I cant do it) x+y+z=4 4x+5y=3 y-3z=-10

OpenStudy (jdoe0001):

have you covered solving "system of equations of two variables" yet?

OpenStudy (anonymous):

yea I have just been working on this for too long and I cant find an answer

OpenStudy (jdoe0001):

k

OpenStudy (jdoe0001):

gimme a few

OpenStudy (jdoe0001):

\(\large { \begin{array}{llll} x+y+z&=4\\ 4x+5y&=3\\ y-3z&=-10 \end{array}\implies \begin{array}{cllll} x+y+z&=4\\ 4x+5y+{\color{brown}{ 0z}}&=3\\ {\color{brown}{ 0x}}+y-3z&=-10 \end{array} \\ \quad \\ \textit{now let us pick the 1st two only} \\ \quad \\ \begin{array}{cllcll} x+y+z&=4&{\color{red}{ \times -4}}\implies &\cancel{ -4x}-4y-4z=-16\\ 4x+5y+0z&=3&&\cancel{ 4x}+5y+0z=3\\ \hline\\ &&&\square ?+\square ?+\square ?=\square ? \end{array} }\) what would that give us as resulting equation?

OpenStudy (anonymous):

9y-4z=-13??

OpenStudy (jdoe0001):

-4y+5y = 9y? *cough*

OpenStudy (jdoe0001):

but anyhow, the others are ok, so \(\large \begin{array}{cllcll} x+y+z&=4&{\color{red}{ \times -4}}\implies &\cancel{ -4x}-4y-4z=-16\\ 4x+5y+0z&=3&&\cancel{ 4x}+5y+0z=3\\ \hline\\ &&&0x+y-4z=-13 \end{array}\) so if we solve there for "y" what would that give us?

OpenStudy (jdoe0001):

0x+y-4z = -13 or y-4z = -13 solve for "y" would give us?

OpenStudy (anonymous):

it would give us a decimal right?

OpenStudy (jdoe0001):

nope

OpenStudy (jdoe0001):

it'd give us a value, in "z' terms

OpenStudy (anonymous):

i have no clue i hate math

OpenStudy (jdoe0001):

hehe well.... you'd need to solve it like you'd any other linear simplification say let us use "a" and "b" if you had a - 4b = 3 solving for "a" would give you?

OpenStudy (jdoe0001):

a - 4b = -13 rather what would you get for "a"?

OpenStudy (jdoe0001):

hint: add 4b to both sides

OpenStudy (anonymous):

a=-13=4b

OpenStudy (jdoe0001):

well... if you add +4b to both sides you'd get ?

OpenStudy (anonymous):

a=13+4b

OpenStudy (anonymous):

a=-13+4b

OpenStudy (jdoe0001):

well....yes.. it just that is -13 though

OpenStudy (anonymous):

yea sorry

OpenStudy (anonymous):

which system was that apart of?

OpenStudy (jdoe0001):

one sec

OpenStudy (jdoe0001):

ohh wel... I simply changed the letters if a -4b = -13 => a = -13 + 4b then y - 4z = -13 => y = -13+4z

OpenStudy (anonymous):

But how does that help solve the system, if i now have y equaled to something?

OpenStudy (jdoe0001):

\(\large { \begin{array}{llll} (1)x+y+z&=4\\ (2)4x+5y&=3\\ (3){\color{blue}{ y}}-3z&=-10 \end{array}\implies \begin{array}{cllll} x+y+z&=4\\ 4x+5y+{\color{brown}{ 0z}}&=3\\ {\color{brown}{ 0x}}+y-3z&=-10 \end{array} \\ \quad \\ y -4z = -13 \implies {\color{blue}{ y}}=\underline{ -13+4z} \\ \quad \\ thus\qquad (3){\color{blue}{ y}}-3z=-10\implies {\color{blue}{ -13+4z}}-3z=-10 }\) can you tell what "z" is now?

OpenStudy (jdoe0001):

notice above since we know now that y =-13+4z then we could use that in the (3) equation and "substitute" "y" for that to get "z"

OpenStudy (anonymous):

oh i see :D

OpenStudy (jdoe0001):

-13+4z-3z = -10 1z -13+z = -10 z = -10 + 13

OpenStudy (jdoe0001):

can you seee what "z" is now?

OpenStudy (anonymous):

z is 3

OpenStudy (jdoe0001):

yeap z = 3 and y = -13+4z or y = -13 + 4(3)

OpenStudy (jdoe0001):

once you have "y" and "z" you can just plug them in the (1) equation and get "x"

OpenStudy (anonymous):

when i check two of them dont work

OpenStudy (jdoe0001):

hmm what values did you get for all 3 variables?

OpenStudy (anonymous):

x=-3 y=4 z=3

OpenStudy (jdoe0001):

y = -13+4z or y = -13 + 4(3) => y =?

OpenStudy (anonymous):

Thank you so much

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

@jdoe0001 what were your variable answers btw?

OpenStudy (jdoe0001):

ahemm well y = -13+4z or y = -13 + 4(3) => y =-13 + 4*3 => y = -13 + 12 y = -1 we know z = 3 and y = -1 plug them in the 1st equation then \(\bf x+{\color{brown}{ y}}+{\color{brown}{ z}}=4\implies x+{\color{brown}{ (-1)}}+{\color{brown}{ (3)}}=4\implies x-1+3=4 \\ \quad \\ x+2=4\implies x\cancel{ +2-2}=4-2\implies x=2\)